# Learning to Orchestrate Agents in Natural Language with the Conductor

### v1

@@ A.2 Datasets
We select the following tasks for training and evaluation both for their usage in other multi-agent works and for their difficulty and popularity in measuring frontier model performance. For instance, MATH500 is used in (yue2025masrouter) and MMLU is used in (yue2025masrouter; chen2024routerdc). We also note that we chose harder versions of popular benchmarks in multi-agent papers. For example, yue2025masrouter; chen2024routerdc; multi_agent_debate use GSM8K to evaluate mathematical reasoning, and we chose the more challenging AIME25.
MMLU (hendrycks2020measuring). A massive multitask language comprehension dataset spanning 57 tasks, including history, law, sciences, and social sciences, among others, testing world knowledge and problem solving ability. All questions are multiple choice. The training data comprises 99842 samples and the test data comprises 14042 samples.
MATH500 (lightman2023lets). A subset of 500 of the most challenging problems selected by lightman2023lets from the MATH dataset (hendrycks2021measuring). We construct a train set of 300 samples and a test set of 100 samples.
LiveCodeBench (jain2024livecodebench). This is the coding subset of the Live Bench (white2024livebench), which is a contamination-free, continuously updated benchmark spanning 17 diverse tasks including coding, math, data analysis, and language among others. LiveCodeBench comprises leetcode-style code generation problems. We select the oldest version, version 1, as our training data, and the newest version (at the time of writing), version 6, for testing.
RLPR (yu2025rlpr). A non-mathematical, general reasoning dataset obtained as a subset from WebInstruct (yue2024mammoth2), where samples are filtered using GPT-4.1 to remove overly easy questions. The total dataset comprises 77700 samples, of which we take 46620 as training data and 15540 as test data.
AIME25 (maa_aime2025). One of the most challenging and popular mathematical reasoning benchmarks, this dataset is the 30 questions used in the 2025 edition of the annual American Invitational Mathematics Examination. We use all 30 questions for evaluation.
GPQA-Diamond (rein2024gpqa). This is the set of diamond difficulty problems on natural science taken from the Graduate-level Google-proof Q&A benchmark. The dataset comprises 198 problems in multiple-choice format. We use all 198 for evaluation.
BigCodeBench (zhuo2024bigcodebench). A challenging code generation benchmark focusing on diverse function calls and challenging instructions. We select the ”hard” and ”complete” subset, comprising 148 samples, which is specifically designed for code completion based on comprehensive docstrings. We use all 148 samples for evaluation. For all code evaluation we use the default gradio backend, hosted on the HuggingFace space [^3] https://huggingface.co/spaces/bigcode/bigcodebench-evaluator.
@@ A.3 Evaluation Setup
Unconstrained setting. We refer to the setting in which we set the completion tokens and reasoning budget limits to their max as unconstrained. This is the setting with which we obtain our results in Section 4.2. Completion tokens and reasoning budgets are detailed in Table 3. In the case of Qwen3-32B (thinking), no budget is configurable, instead we simply select ”enabled”. We make one exception to the max reasoning configuration in the unconstrained evaluation setup, which is for GPT-5 in BigCodeBench, where we evaluate GPT-5 at medium reasoning effort as opposed to high reasoning effort. This is due to GPT-5’s performance being marginally stronger under medium reasoning effort, which is in line with OpenAI’s own findings regarding GPT-5 (openai2025introducing-gpt5), in which for certain tasks medium may outperform high reasoning effort.
@@
We evaluate AIME25 and GPQA-Diamond using Lighteval (fourrierlighteval) and BigCodeBench using the original source repository [^4] https://github.com/bigcode-project/bigcodebench.
Baselines. Within the constrained setting, we evaluate against state-of-the-art multi-agent routing and scaffolding techniques, including MasRouter (yue2025masrouter), RouterDC (chen2024routerdc), Smoothie (guha2024smoothie), and MoA (wang2024mixture) as seen in Figure 4 and Table 7. We train MasRouter and RouterDC using the by sampling from the training dataset that was used to train Conductor. We select the evaluation model based on best validation loss from a validation set which is of the same distribution, but independent of both the training and testing sets. Specifically, MasRouter is trained with a batch size 256 and validated every 5 iterations. We stop training early after seeing sufficient evidence of overfitting (i.e. $10\%$ drop off in validation). RouterDC is trained on 500 samples, with each sample repeated 5 times to collect an average performance of every work on the given question. The collected average performance is used to compute the contrastive loss described in chen2024routerdc. Smoothie is applied to test-time questions and outputs of each worker model with both dependent (selecting a model per question) and independent (selecting a single model for all questions) strategies. MoA is also applied as a test-time scaffold with a single MoA layer and single aggregator layer for a total of 8 model calls. The aggregator model is chosen at random. All baselines were compared against using the code and default settings provided by their respective authors.
@@
Performance improvement scale. We note that in our evaluation setup we focus on highly competitive reasoning benchmarks, such as AIME and GPQA-Diamond. Such tasks tend to have a long-tailed distribution of difficulty (xu2025unveiling) where breakthroughs in a small subset of particularly challenging problems could be representative of entire generational improvements in LLM reasoning. For example, the difference between two generations of GPT reasoning models, from GPT-o3 and GPT-5, is 3.3% on AIME25555Kaggle leadboard: https://www.kaggle.com/benchmarks/open-benchmarks/aime-2025 and 2.7% on GPQA-Diamond [^6] Artificial Analysis leaderboard: https://shorturl.at/eLHUj in absolute percentage terms. Moreover, the Conductor framework not only yields improvements in a single domain but across a diverse range of such highly challenging benchmarks, across math, coding, and natural science, something we believe particularly validates the meaningfulness of its advancements beyond the performance foundation of frontier models.
@@ B.1 Efficiency Analysis
We present in Table 5 additional efficiency results for the Conductor in comparison with Claude Sonnet 4, Gemini 2.5 Pro, and GPT-5 using a consensus inference-time scaling framework (wang2022self). We take MMLU as a representative task, set the consensus sampling to 5 (in keeping with the Conductor’s max allowable workflows of 5 steps), and report average token usage per sample, average cost per sample, and cost adjusted performance (taken simply as performance / average cost per sample in cents). We see the Conductor not only outperforms consensus, but additionally offers substantial efficiency gains in terms of token usage and average cost relative to this popular inference-time scaling technique.
@@ B.3 Performance Diversity
Throughout our evaluation, we find, in line with existing leaderboards and surveys (Chang2024LLMEvalSurvey), that no single model reigns supreme over all tasks, with differing models excelling or struggling in differing tasks. Examples of this in our own evaluation include GPT-5’s strong performance in math and competitive coding (seen in AIME and LiveCodeBench), while Gemini excels in scientific reasoning (GPQA-Diamond). Claude Sonnet 4 struggles at competitive coding (relatively weak in LiveCodeBench), but is one of the dominant models at code generation with diverse function calling (BigCodeBench). This specialization is also prevalent at the granular ”sub-task” level, where we often find the Conductor learns that different models are most useful as ”planners” or ”writers” to answer particular kinds of questions. For instance, our SOTA performance in LiveCodeBench leverages Gemini 2.5 Pro and Claude Sonnet 4 working together as high-level planners and only later employs GPT-5 to write the final optimized code, which far outperformed alternate strategies using GPT-5 in a planning role that the Conductor attempted at early training iterations.
@@ Appendix E Conductor Prompt and Few-Shot Examples
Few-shot example details. As detailed in Section 4, we design our training set as a combination of problems taken from MATH500, LiveCodeBench, MMLU, and RLPR. Our OOD few-shot examples are taken from Medreason (wu2025medreasonelicitingfactualmedical), DeepMath (deepmath), and Countdown (pan2025learning). We select just four examples, one from each of Medreason and DeepMath and two from Countdown, and select examples to ensure a range of workflow steps and selected agents to encourage exploration of the coordination space. Examples shown in Figure 17. For the In-Distribution few-shot setup, we again choose four few-shot examples, with one from each of our component tasks, MATH500, RLPR, MMLU, and LiveCodeBench. Again, we select examples to balance workflow steps and agent selection to encourage exploration. Examples shown in Figure 16. For the semi-OOD setting, we balance in-distribution and OOD, taking one example from MATH500, one example from LiveCodeBench, one example from Countdown, and one example from Medreason.

### v2

@@ A.2 Datasets
We select the following tasks for training and evaluation both for their usage in other multi-agent works and for their difficulty and popularity in measuring frontier model performance. For instance, MATH500 is used in (Yue et al., 2025) and MMLU is used in (Yue et al., 2025; Chen et al., 2024). We also note that we chose harder versions of popular benchmarks in multi-agent papers. For example, Yue et al. (2025); Chen et al. (2024); Du et al. (2023b) use GSM8K to evaluate mathematical reasoning, and we chose the more challenging AIME25.
MMLU (Hendrycks et al., 2020). A massive multitask language comprehension dataset spanning 57 tasks, including history, law, sciences, and social sciences, among others, testing world knowledge and problem solving ability. All questions are multiple choice. The training data comprises 99842 samples and the test data comprises 14042 samples.
MATH500 (Lightman et al., 2023). A subset of 500 of the most challenging problems selected by Lightman et al. (2023) from the MATH dataset (Hendrycks et al., 2021). We construct a train set of 300 samples and a test set of 100 samples.
LiveCodeBench (Jain et al., 2024). This is the coding subset of the Live Bench (White et al., 2024), which is a contamination-free, continuously updated benchmark spanning 17 diverse tasks including coding, math, data analysis, and language among others. LiveCodeBench comprises leetcode-style code generation problems. We select the oldest version, version 1, as our training data, and the newest version (at the time of writing), version 6, for testing.
RLPR (Yu et al., 2025). A non-mathematical, general reasoning dataset obtained as a subset from WebInstruct (Yue et al., 2024), where samples are filtered using GPT-4.1 to remove overly easy questions. The total dataset comprises 77700 samples, of which we take 46620 as training data and 15540 as test data.
AIME25 (Mathematical Association of America, 2025). One of the most challenging and popular mathematical reasoning benchmarks, this dataset is the 30 questions used in the 2025 edition of the annual American Invitational Mathematics Examination. We use all 30 questions for evaluation.
GPQA-Diamond (Rein et al., 2024). This is the set of diamond difficulty problems on natural science taken from the Graduate-level Google-proof Q&A benchmark. The dataset comprises 198 problems in multiple-choice format. We use all 198 for evaluation.
BigCodeBench (Zhuo et al., 2024). A challenging code generation benchmark focusing on diverse function calls and challenging instructions. We select the ”hard” and ”complete” subset, comprising 148 samples, which is specifically designed for code completion based on comprehensive docstrings. We use all 148 samples for evaluation. For all code evaluation we use the default gradio backend, hosted on the HuggingFace space [^3] https://huggingface.co/spaces/bigcode/bigcodebench-evaluator.
@@ A.3 Evaluation Setup
Unconstrained setting. We refer to the setting in which we set the completion tokens and reasoning budget limits to their max as unconstrained. This is the setting with which we obtain our results in Section 4.2. Completion tokens and reasoning budgets are detailed in Table 3. In the case of Qwen3-32B (thinking), no budget is configurable, instead we simply select ”enabled”. We make one exception to the max reasoning configuration in the unconstrained evaluation setup, which is for GPT-5 in BigCodeBench, where we evaluate GPT-5 at medium reasoning effort as opposed to high reasoning effort. This is due to GPT-5’s performance being marginally stronger under medium reasoning effort, which is in line with OpenAI’s own findings regarding GPT-5 (OpenAI, 2025), in which for certain tasks medium may outperform high reasoning effort.
@@
We evaluate AIME25 and GPQA-Diamond using Lighteval (Fourrier et al., ) and BigCodeBench using the original source repository [^4] https://github.com/bigcode-project/bigcodebench.
Baselines. Within the constrained setting, we evaluate against state-of-the-art multi-agent routing and scaffolding techniques, including MasRouter (Yue et al., 2025), RouterDC (Chen et al., 2024), Smoothie (Guha et al., 2024), and MoA (Wang et al., 2024) as seen in Figure 4 and Table 7. We train MasRouter and RouterDC using the by sampling from the training dataset that was used to train Conductor. We select the evaluation model based on best validation loss from a validation set which is of the same distribution, but independent of both the training and testing sets. Specifically, MasRouter is trained with a batch size 256 and validated every 5 iterations. We stop training early after seeing sufficient evidence of overfitting (i.e. $10\%$ drop off in validation). RouterDC is trained on 500 samples, with each sample repeated 5 times to collect an average performance of every work on the given question. The collected average performance is used to compute the contrastive loss described in Chen et al. (2024). Smoothie is applied to test-time questions and outputs of each worker model with both dependent (selecting a model per question) and independent (selecting a single model for all questions) strategies. MoA is also applied as a test-time scaffold with a single MoA layer and single aggregator layer for a total of 8 model calls. The aggregator model is chosen at random. All baselines were compared against using the code and default settings provided by their respective authors.
@@
Performance improvement scale. We note that in our evaluation setup we focus on highly competitive reasoning benchmarks, such as AIME and GPQA-Diamond. Such tasks tend to have a long-tailed distribution of difficulty (Xu et al., 2025) where breakthroughs in a small subset of particularly challenging problems could be representative of entire generational improvements in LLM reasoning. For example, the difference between two generations of GPT reasoning models, from GPT-o3 and GPT-5, is 3.3% on AIME25555Kaggle leadboard: https://www.kaggle.com/benchmarks/open-benchmarks/aime-2025 and 2.7% on GPQA-Diamond [^6] Artificial Analysis leaderboard: https://shorturl.at/eLHUj in absolute percentage terms. Moreover, the Conductor framework not only yields improvements in a single domain but across a diverse range of such highly challenging benchmarks, across math, coding, and natural science, something we believe particularly validates the meaningfulness of its advancements beyond the performance foundation of frontier models.
@@ B.1 Efficiency Analysis
We present in Table 5 additional efficiency results for the Conductor in comparison with Claude Sonnet 4, Gemini 2.5 Pro, and GPT-5 using a consensus inference-time scaling framework (Wang et al., 2022). We take MMLU as a representative task, set the consensus sampling to 5 (in keeping with the Conductor’s max allowable workflows of 5 steps), and report average token usage per sample, average cost per sample, and cost adjusted performance (taken simply as performance / average cost per sample in cents). We see the Conductor not only outperforms consensus, but additionally offers substantial efficiency gains in terms of token usage and average cost relative to this popular inference-time scaling technique.
@@ B.3 Performance Diversity
Throughout our evaluation, we find, in line with existing leaderboards and surveys (Chang et al., 2024), that no single model reigns supreme over all tasks, with differing models excelling or struggling in differing tasks. Examples of this in our own evaluation include GPT-5’s strong performance in math and competitive coding (seen in AIME and LiveCodeBench), while Gemini excels in scientific reasoning (GPQA-Diamond). Claude Sonnet 4 struggles at competitive coding (relatively weak in LiveCodeBench), but is one of the dominant models at code generation with diverse function calling (BigCodeBench). This specialization is also prevalent at the granular ”sub-task” level, where we often find the Conductor learns that different models are most useful as ”planners” or ”writers” to answer particular kinds of questions. For instance, our SOTA performance in LiveCodeBench leverages Gemini 2.5 Pro and Claude Sonnet 4 working together as high-level planners and only later employs GPT-5 to write the final optimized code, which far outperformed alternate strategies using GPT-5 in a planning role that the Conductor attempted at early training iterations.
@@ Appendix E Conductor Prompt and Few-Shot Examples
Few-shot example details. As detailed in Section 4, we design our training set as a combination of problems taken from MATH500, LiveCodeBench, MMLU, and RLPR. Our OOD few-shot examples are taken from Medreason (Wu et al., 2025), DeepMath (He et al., 2025), and Countdown (Pan et al., 2025). We select just four examples, one from each of Medreason and DeepMath and two from Countdown, and select examples to ensure a range of workflow steps and selected agents to encourage exploration of the coordination space. Examples shown in Figure 17. For the In-Distribution few-shot setup, we again choose four few-shot examples, with one from each of our component tasks, MATH500, RLPR, MMLU, and LiveCodeBench. Again, we select examples to balance workflow steps and agent selection to encourage exploration. Examples shown in Figure 16. For the semi-OOD setting, we balance in-distribution and OOD, taking one example from MATH500, one example from LiveCodeBench, one example from Countdown, and one example from Medreason.